# A Training-Free Style-aligned Image Generation with Scale-wise Autoregressive Model

### v1

@@ 3.5 Dynamic style injection
where the updated $\mathbf{V}_{s}$ serves as the input to the cross-attention module $M_{CA}(\cdot)$. The interpolation weight $\alpha_{s}$ is dynamically adjusted using a non-linear decreasing schedule function $f_{sc}$, inspired by the concave increasing trend observed during the mid-stage (steps 3–7) of the style-content similarity graph in Fig. 3, which follows an overall S-curve pattern. This function promotes stronger style injection during the mid-stage, while gradually reducing its influence in later steps to allow for natural content refinement. The schedule function is defined as:
@@
Metric Ours StyleAligned [19] B-LoRA [13] StyleDrop [44] DB-LoRA [40] IP-Adapter [56]
Object relevancy ($S_{\text{obj}}$) $\uparrow$ 0.282 0.281 0.302 0.267 0.309 0.277
Style consistency ($S_{\text{sty}}$) $\uparrow$ 0.551 0.530 0.292 0.425 0.369 0.529
Dual Consistency ($S_{\text{dual}}$) $\uparrow$ 0.373 0.367 0.270 0.328 0.336 0.364
Inference time (seconds) $\downarrow$ 1.58 11.25 633.20 544.06 343.68 10.14
Table 1: Quantitative comparison with state-of-the-art style-aligned image generation models. We evaluate the generated image sets in terms of object relevancy (CLIP score), style consistency (DINO embedding similarity), and dual consistency (harmonic mean of the CLIP score and DINO embedding similarity). The inference time is measured per image. The symbols $\uparrow$ and $\downarrow$ indicate higher values are better and lower values are better, respectively.
@@ 4.1 Implementation Details
We implement our method using pre-trained Infinity [17] 2B model with all parameters frozen, which performs scale-wise prediction across 12 steps. Initial feature replacement is applied at all early steps $\mathbf{S}_{\text{e}}$ (see Sec. 3.2), followed by pivotal feature interpolation at the 3rd scale ($\bar{s}=3$) with a fixed interpolation weight of $\alpha_{\text{pivot}}=0.4$. Alternatively, $\bar{s}$ can also be applied at any mid-stage scale in $\mathbf{S}_{\text{m}}$ as a hyperparameter, with minimal performance variation (see supplementary material), demonstrating the robustness of our model. Additionally, dynamic style injection is applied at all mid-stage $\mathbf{S}_{\text{m}}$, where the interpolation weight $\alpha_{s}$ is determined by a schedule function with decay rate $r$ set to 3.4. Our baseline uses a codebook with a dimension of $2^{32}$, and the quantized feature map has a resolution of 64$\times$64$\times$32. The number of generated style-aligned images can be controlled by specifying the number of input text prompts. Generating a set of four 1024$\times$1024 images simultaneously takes approximately 6.3 seconds (1.58 seconds per image) on a single A6000 GPU.
@@ 4.2 Evaluation metrics
Following [19], we adopt the same evaluation methodology to assess our method based on two key aspects: object relevancy ($S_{\text{obj}}$) and style consistency ($S_{\text{sty}}$) across the generated image sets. To evaluate object relevancy, we compute the CLIP cosine similarity [36] between each generated image and its corresponding text description, ensuring that the generated content aligns with the given prompt. For style consistency, we measure the pairwise average cosine similarity of DINO ViT-B/8 [3] embeddings within each generated set. Consistent with [19], we use DINO embeddings instead of CLIP image embeddings, as CLIP, trained with class labels, tends to produce high similarity scores for images with similar content even if they have differing styles. In contrast, DINO embeddings, trained in a self-supervised manner, are more sensitive to style variations, making them more suitable for evaluating style alignment. Additionally, we propose a combined metric, dual consistency ($S_{\text{dual}}$) to provide a balanced assessment of both object relevancy and style consistency. We compute it as the harmonic mean of the two metrics $S_{\text{dual}}=2S_{\text{obj}}S_{\text{sty}}/\big{(}{S_{\text{obj}}+S_{\text{sty}}}\big{)}$. For a fair comparison, we use the same set of 100 text prompts generated by ChatGPT, as introduced in [19].
@@
# Re PFI DSI $S_{\text{dual}}$uparrow $S_{\text{obj}}$uparrow $S_{\text{sty}}$uparrow
@@
(d) 0.373 0.282 0.551
Table 2: Ablation study on the initial feature replacement (Re), Pivotal Feature Interpolation (PFI), Dynamic style injection (DSI). The symbol $\uparrow$ indicates higher values are better, respectively.
Quantitative comparison. Tab. 1 presents a quantitative comparison between our method and state-of-the-art style-aligned image generation models: StyleAligned [19], B-LoRA [13], StyleDrop [44], DreamBooth-LoRA (DB-LoRA) [40], and IP-Adapter [56]. The results demonstrate that our method achieves the best balance between object relevancy and style consistency, while also significantly outperforming all compared models in inference speed. Our method surpasses StyleAligned [19] across all metrics, while being approximately 6×6$\times$6 × faster. Although our object relevancy score is slightly lower than B-LoRA [13] and DB-LoRA [40], our method achieves substantially higher style consistency, leading to the highest score in the dual consistency metric. This demonstrates the effectiveness of our approach in harmonizing object relevancy and style consistency to ensure high-quality, style-aligned image generation. Furthermore, our method is approximately 400×400$\times$400 × faster than B-LoRA [13] and 200×200$\times$200 × faster than DB-LoRA [40], highlighting its exceptional efficiency.
Qualitative comparison. To further demonstrate the effectiveness of our proposed method, we present a qualitative comparison with existing models in Fig. 5. Compared to B-LoRA [13] and DB-LoRA [40], which achieve strong object fidelity due to their high object relevancy scores, both methods struggle with maintaining consistent style across all images in a set. While the generated images accurately depict the specified objects, they often fail to consistently apply the desired artistic attributes, resulting in a lack of cohesive visual appearance. StyleAligned [19] demonstrates relatively stronger style consistency than B-LoRA [13] and DB-LoRA [40] and is the best-performing among existing models. However, it still falls short of our approach, occasionally generating repetitive patterns or artifacts that reduce object diversity and compromise visual quality. In contrast, our method successfully integrates both object relevancy and style consistency. For more challenging styles, such as glowing and paper cutout styles, our model effectively captures the intended stylistic characteristics while preserving clear, well-formed object structures. These qualitative results validate the effectiveness of our approach in achieving high-quality, style-aligned image generation.
@@ 4.4 Ablation study
Quantitative analysis. The quantitative results in Tab. 2 validate the contributions of each proposed technique to improving both style consistency and overall image quality. Each component offers distinct benefits. As shown in Tab. 2-(b), Initial feature replacement enhances style consistency by ensuring a uniform RGB statistics, leading to a clear increase in $S_{\text{sty}}$. In Tab. 2-(c), Pivotal feature interpolation further improves style alignment while preserving object relevancy. Integrating all components including dynamic style injection, as shown in Tab. 2-(d), significantly boosts style consistency results while resulting in a slight decrease in object relevancy. However, this minor-trade off is outweighed by substantial gains in style consistency and dual consistency. These results confirm that each module effectively contributes to improving style alignment while preserving content fidelity across the generated image sets.
Qualitative analysis. The qualitative results in Fig. 6 demonstrate the impact of applying or omitting each proposed method on the generated images. As seen in Fig. 6-(a), the absence of our proposed techniques results in misaligned styles across images, with noticeable variations in color tones and lighting conditions. Fig. 6-(b), incorporating Initial feature replacement, produces more uniform background colors, improving overall style consistency while maintaining content fidelity. In Fig. 6-(c), where Pivotal feature interpolation is additionally applied, the generated images exhibit enhanced coherence in object placement and style alignment. Finally, Fig. 6-(d) integrates all three proposed techniques, including Dynamic style injection, which further refines style consistency while preserving content structure. These qualitative ablation results demonstrate that each module is crucial in improving style alignment while maintaining content fidelity across the generated image sets.
@@
function $S_{\text{dual}}$uparrow $S_{\text{obj}}$uparrow $S_{\text{sty}}$uparrow
@@
Ours $f_{sc}(s)$ 0.373 0.282 0.551
Table 3: Additional ablation study on various schedule functions.
Figure 7: Visualization of various schedule functions.
@@ 4.5 Schedule function design
We conduct an additional ablation study to evaluate the effectiveness of our schedule function. Based on our observations in Fig. 3, we hypothesize that gradually incorporating style information during the stages where style and content are progressively defined would enhance style alignment while preserving the intended content structure. To test this hypothesis, we design five different schedule functions (illustrated in Fig. 7), each gradually decreasing over the generation steps, to assess their ability to promote style consistency without compromising content fidelity.
As shown in Tab. 3, our proposed schedule function – derived directly from our observations – achieves the highest $S_{\text{dual}}$ score, demonstrating its effectiveness in balancing both object relevancy and style consistency. Notably, the concave up schedule, which follows a decay pattern similar to our design, also performs relatively well, further supporting our hypothesis. In contrast, the linear, constant, concave-down, and cosine schedules, which deviate from our observed trends, exhibit lower performance. These results confirm that our observation-driven schedule function optimally integrates style while preserving content, highlighting the superiority of our approach.
4.6 User study
We conduct a user study to further enhance our evaluation, with the results presented in Tab. 4. The study involved 50 participants, ranging in age from their 20s to 50s. Participants were asked to evaluate two key aspects: object relevancy and style consistency. We compare images generated by our model with those produced by StyleAligned [19], StyleDrop [44], and IP-Adapter [56], the top three performers in quantitative evaluations for style-aligned image generation. The user study results demonstrate that our model effectively preserves object relevancy while achieving strong style consistency across the generated image sets.
Method Object $\uparrow$ Style $\uparrow$
StyleAligned [19] 14.2 % 31.5 %
StyleDrop [44] 3.3 % 2.2 %
IP-Adapter [56] 27.0% 13.5 %
Ours 55.5 % 52.8 %
Table 4: User study detailing preference percentages.
@@ 5 Conclusion
In this paper, we introduce an efficient and practical training-free style-aligned image generation method built on a scale-wise autoregressive model. To achieve this, we analyze the next-scale prediction scheme in terms of color statistics, content evolution, and style variations across generation steps. We observe that the early-stage primarily establishes the overall RGB statistics of the generated images, setting the foundation for the final appearance. In the mid-stage, content and style progressively evolve, incorporating finer details and refining visual coherence.
Building on these insights, we propose three key components to enhance style alignment and content preservation. Initial feature replacement ensures consistent RGB statistics by controlling early-stage features, which heavily influence overall color distribution. Pivotal feature interpolation guides object placement and visual coherence at one of the critical middle steps. Finally, Dynamic style injection gradually refines self-attention interpolation using a schedule function, balancing style consistency with content preservation.
Extensive experiments demonstrate that our method achieves the highest dual consistency while ensuring the fastest inference time, delivering high-quality, style-consistent generations without additional fine-tuning or training. We hope this work inspires further research into training-free approaches that enhance the controllability and efficiency of autoregressive T2I models.
@@
6 Comprehensive Analysis of our method
6.1 Additional analysis on generation process
To verify whether our method operates as intended during the generation process, we conduct an additional analysis at each step using 400 images generated from our method and the baseline model, Infinity [17]. Specifically, we compare the first image in each batch with the subsequent images at different generation steps to examine how RGB statistics, style, and content evolve over time.

### v2

@@ 3.5 Dynamic style injection
where the updated $\mathbf{V}_{s}$ serves as the input to the self-attention module $M_{SA}(\cdot)$.
The interpolation weight $\alpha_{s}$ is dynamically adjusted using a non-linear decreasing schedule function $f_{sc}$, inspired by the concave increasing trend observed during the mid-stage (steps 3–7) of the style-content similarity graph in Fig. 3, which follows an overall S-curve pattern. This design encourages stronger style injection at the beginning of the mid-stage—when global style attributes are still forming—and gradually weakens it toward the later steps, allowing content features to stabilize and refine naturally. The schedule function is defined as:
@@
Table 1: Quantitative comparison with state-of-the-art style-aligned image generation models. We evaluate the generated image sets in terms of object relevancy (CLIP text score), style consistency (DINO embedding similarity), and dual consistency (harmonic mean of the CLIP score and DINO embedding similarity). The inference time is measured per image, where for DB-LoRA, B-LoRA, and StyleDrop, the reported time includes both inference and per-style optimization. The symbols $\uparrow$ and $\downarrow$ indicate that higher values are better and lower values are better, respectively.
Metric Ours StyleAligned B-LoRA StyleDrop DB-LoRA IP-Adapter CSGO AlignedGen DreamO
Training-free
Object relevancy ($S_{\text{obj}}$) $\uparrow$ 0.282 0.281 0.302 0.267 0.309 0.277 0.292 0.283 0.292
Style consistency ($S_{\text{sty}}$) $\uparrow$ 0.556 0.530 0.292 0.425 0.369 0.529 0.432 0.486 0.432
Dual Consistency ($S_{\text{dual}}$) $\uparrow$ 0.375 0.367 0.270 0.328 0.336 0.364 0.348 0.358 0.348
Inference time (seconds) $\downarrow$ 1.58 11.25 633.20 544.06 343.68 10.14 14.80 45.21 14.95
@@ 4.1 Implementation Details
We implement our method using pre-trained Infinity [han2024infinity] 2B model with all parameters frozen, which performs scale-wise prediction across 12 steps. Initial feature replacement is applied at all early steps $\mathbf{S}_{\text{e}}$ (see Sec. 3.2), followed by pivotal feature interpolation at the 3rd scale ($\bar{s}=3$) with a fixed interpolation weight of $\alpha_{\text{pivot}}=0.4$. Alternatively, $\bar{s}$ can also be applied at any mid-stage scale in $\mathbf{S}_{\text{m}}$ as a hyperparameter, with minimal performance variation (see supplementary material), demonstrating the robustness of our model. Additionally, dynamic style injection is applied at all mid-stage $\mathbf{S}_{\text{m}}$, where the interpolation weight $\alpha_{s}$ is determined by a schedule function with decay rate $r$ set to 3.4. Our baseline uses a codebook with a dimension of $2^{32}$, and the quantized feature map has a resolution of $\times64$$\times$32. The number of generated style-aligned images can be controlled by specifying the number of input text prompts. Generating a set of four $\times1024$ images simultaneously takes approximately 6.3 seconds (1.58 seconds per image) on a single A6000 GPU.
@@ 4.2 Evaluation metrics
Following [hertz2024stylealignedimagegeneration], we adopt the same evaluation methodology to assess the generated image sets based on two key aspects: object relevancy ($S_{\text{obj}}$) and style consistency ($S_{\text{sty}}$). To evaluate object relevancy, we compute the CLIP cosine similarity [radford2021learning] between each generated image and its corresponding text description, ensuring that the generated content aligns with the given prompt. To evaluate style consistency, we utilize DINO ViT-B/8 [caron2021emerging], which is sensitive to style variations, following the approach of [hertz2024stylealignedimagegeneration]. We then measure the pairwise average cosine similarity of the DINO embeddings within each generated set. Additionally, we propose a combined metric, dual consistency ($S_{\text{dual}}$), to provide a balanced assessment of both object relevancy and style consistency. We compute it as the harmonic mean of the two metrics $S_{\text{dual}}=2S_{\text{obj}}S_{\text{sty}}/\big({S_{\text{obj}}+S_{\text{sty}}}\big)$. For a fair comparison, we adopt the same 100 text prompts generated by ChatGPT as used in [hertz2024stylealignedimagegeneration], generating 4 image sets per prompt and totaling 400 images for evaluation.
Quantitative comparison. Tab. 1 presents a quantitative comparison between our method and state-of-the-art style-aligned image generation models: StyleAligned [hertz2024stylealignedimagegeneration], B-LoRA [frenkel2024implicitstylecontentseparationusing], StyleDrop [sohn2023styledrop], DreamBooth-LoRA (DB-LoRA) [ryu2023low], IP-Adapter [ye2023ipadaptertextcompatibleimage], CSGO [xing2024csgo], AlignedGen [zhang2025alignedgen], and DreamO [mou2025dreamo]. The results demonstrate that our method achieves the best balance between object relevancy and style consistency, while also significantly outperforming all compared models in inference speed.
[…]
Qualitative comparison. To further demonstrate the effectiveness of our proposed method, we present a qualitative comparison with existing models in Fig. 5. B-LoRA [frenkel2024implicitstylecontentseparationusing] and DB-LoRA [ryu2023low] exhibit strong object fidelity owing to their high object relevancy, yet their results reveal a clear lack of style consistency compared to our model and other training-free approaches. StyleAligned [hertz2024stylealignedimagegeneration] and AlignedGen [zhang2025alignedgen] produce results visually comparable to ours, but subtle differences remain: in StyleAligned, the generated objects are occasionally incomplete or distorted (e.g., in the first and second rows), while AlignedGen shows relatively inconsistent style coherence across images (e.g., in the second and third rows). In contrast, the results show that our method consistently preserves both object structure and stylistic uniformity. These qualitative results demonstrate that our approach achieves a more stable balance between object fidelity and style consistency, leading to high-quality, style-aligned generations.
Table 2: Ablation study on the initial feature replacement (Re), Pivotal Feature Interpolation (PFI), Dynamic style injection (DSI). The symbol $\uparrow$ indicates that higher values are better.
@@ 4.3 Comparison with state-of-the-art style-aligned image generation models
# Re PFI DSI $S_{\text{dual}}$ $\uparrow$ $S_{\text{obj}}$ $\uparrow$ $S_{\text{sty}}$ $\uparrow$
@@
(d) 0.375 0.282 0.556
@@ 4.4 Ablation study
Quantitative analysis. The quantitative results in Tab. 2 validate the contributions of each proposed technique to improving both style consistency and overall image quality. Each component offers distinct benefits. As shown in Tab. 2-(b), Initial feature replacement enhances style consistency by ensuring uniform RGB statistics, leading to a clear increase in $S_{\text{sty}}$. In Tab. 2-(c), Pivotal feature interpolation further improves style alignment while preserving object relevancy. Integrating all components including dynamic style injection, as shown in Tab. 2-(d), significantly boosts style consistency results while resulting in a slight decrease in object relevancy. However, this minor trade-off is outweighed by substantial gains in style consistency and dual consistency. These results confirm that each module effectively contributes to improving style alignment while preserving content fidelity across the generated image sets.
Qualitative analysis. The qualitative results also reinforce these findings. As seen in Fig. 6-(a), the absence of our proposed techniques results in misaligned styles across images, with noticeable variations in color tones and lighting conditions. Fig. 6-(b), incorporating Initial feature replacement, produces more uniform background colors, improving overall style consistency. In Fig. 6-(c), where Pivotal feature interpolation is additionally applied, the generated images exhibit enhanced coherence in object placement and style alignment. Finally, Fig. 6-(d) integrates all three proposed techniques, including Dynamic style injection, which further refines style consistency while preserving content structure. These qualitative ablation results demonstrate that each module is crucial in improving style alignment while maintaining object fidelity across the generated image sets.
4.5 User study
We conduct a user study to further enhance our evaluation, with the results presented in Tab. 3. The study involved 50 participants, ranging in age from their 20s to 50s. Participants were asked to evaluate two key aspects: object relevancy and style consistency. We compare images generated by our model with those produced by StyleAligned [hertz2024stylealignedimagegeneration], AlignedGen [zhang2025alignedgen], and IP-Adapter [ye2023ipadaptertextcompatibleimage], the top three performers in quantitative evaluations for style-aligned image generation. The user study results demonstrate that our model effectively preserves object relevancy while achieving strong style consistency across the generated image sets.
Table 3: User study detailing preference percentages.
Method Object $\uparrow$ Style $\uparrow$
StyleAligned [hertz2024stylealignedimagegeneration] 16.9 % 10.9 %
AlignedGen [zhang2025alignedgen] 29.7 % 32.8 %
IP-Adapter [ye2023ipadaptertextcompatibleimage] 8.0% 17.8 %
Ours 45.4 % 38.5 %
Figure 7: Visualization of various schedule functions.
Table 4: Additional ablation study on various schedule functions.
@@ 4.5 User study
function $S_{\text{dual}}$ $\uparrow$ $S_{\text{obj}}$ $\uparrow$ $S_{\text{sty}}$ $\uparrow$
@@
Ours $f_{sc}(s)$ 0.375 0.282 0.556
@@ 4.6 Schedule function design
Based on our observations in Fig. 3, we hypothesize that gradually incorporating style information during the stages where style and content are progressively defined would enhance style alignment while preserving the intended content structure. To validate this hypothesis, we conduct an additional ablation study by designing five distinct schedule functions(illustrated in Fig. 7) that progressively decay over the generation steps and evaluate their impact.
As shown in Tab. 4, our proposed schedule function, derived directly from our observations, achieves the highest $S_{\text{dual}}$ score, demonstrating its effectiveness in balancing both object relevancy and style consistency. Notably, the concave up schedule, which follows a decay pattern similar to our design, also performs relatively well, further supporting our hypothesis. In contrast, the linear, constant, concave-down, and cosine schedules, which deviate from our observed trends, exhibit lower performance. These results confirm that our observation-driven schedule function optimally integrates style while preserving content.
4.7 Incorporating ours to other models
To demonstrate the generalization ability of our method, we applied it to alternative weights of the baseline models, Infinity 8B, and to another T2I model using scale-wise autoregressive generation, Switti [voronov2024swittidesigningscalewisetransformers]. As shown in Tab. 5, our method significantly improves $S_{\text{dual}}$, indicating effective style alignment while maintaining object fidelity. Notably, the autoregressive model Switti demonstrates consistent style alignment despite having a different architecture from our baseline. These results support the generality of our approach across varying model configurations.
Table 5: Incorporating ours to other scale-wise autoregressive-based Text-to-Image models.
Method $S_{\text{dual}}$ $\uparrow$ $S_{\text{obj}}$ $\uparrow$ $S_{\text{sty}}$ $\uparrow$
Switti 0.308 0.294 0.323
Switti + Ours 0.365 0.281 0.521
Infinity 8B 0.316 0.295 0.340
Infinity 8B + Ours 0.373 0.274 0.585
@@ 5 Conclusion
In this paper, we introduce a practical and efficient training-free style-aligned image generation method built on a scale-wise autoregressive model. To this end, we analyze the next-scale prediction scheme in terms of color statistics, content evolution, and style variation across generation steps. We observe that the early-stage primarily determines the overall RGB statistics, forming the foundation for the final appearance, while the mid-stage progressively refines both content and style to achieve visual coherence.
Building on these insights, we propose three components to enhance style alignment and content preservation. Initial feature replacement ensures consistent RGB statistics by stabilizing early-stage features. Pivotal feature interpolation improves object placement and spatial coherence at critical mid-stage steps. Dynamic style injection gradually refines self-attention interpolation via a scheduling function, balancing style consistency with content fidelity. Extensive experiments demonstrate that our method achieves the highest dual consistency with the fastest inference, delivering high-quality, style-consistent generations without additional fine-tuning or training. We hope this work inspires further research that improves the controllability and efficiency of scale-wise autoregressive T2I models.
[…]
@@
Appendix A Evaluation Details
A.1 Evaluation protocol
Following the evaluation protocol established by StyleAligned [hertz2024stylealignedimagegeneration], we evaluate both the object relevancy and style consistency of the generated images. We further introduce a unified metric, dual consistency, to jointly assess these two aspects.
Object relevancy ($S_{\text{obj}}$).
To evaluate object relevancy, we compute the CLIP cosine similarity [radford2021learning] between each generated image and its corresponding object description, measuring how faithfully the image reflects the given object prompt. We calculate the CLIP-T score for every image–prompt pair and report the average across all samples.
Style consistency ($S_{\text{sty}}$).
To assess style alignment among images conditioned on the same identity prompt, we compute the average pairwise DINO similarity, following prior work [hertz2024style, frenkel2024implicitstylecontentseparationusing]. Specifically, we use the CLS-token features from DINO ViT-B/8 [caron2021emerging], which capture global style properties such as texture, rendering, and background.
Dual consistency ($S_{\text{dual}}$).
To jointly capture object relevancy and style consistency, we compute the harmonic mean of the two metrics:
$S_{\text{dual}}=\frac{2S_{\text{obj}}S_{\text{sty}}}{S_{\text{obj}}+S_{\text{sty}}}.$ (8)
Implementation.
We use the official evaluation data provided by StyleAligned [hertz2024stylealignedimagegeneration], with minor extensions to incorporate the dual consistency metric. All models are evaluated under the same hardware setup, using a single NVIDIA A6000 GPU and PyTorch for implementation.
A.2 Implementation details for comparison models
All comparison models are run using the official hyperparameters provided in their public repositories without additional tuning.
Image-based style-aligned image generation models.
We compare our method against several reference image-based style-aligned image generation models that leverage Stable Diffusion XL [podell2023sdxlimprovinglatentdiffusion] or Muse [chang2023musetexttoimagegenerationmasked] as their backbone. Specifically, we include the following representative models. For a fair comparison, we use each model’s first generated output for a given prompt as its reference image, ensuring that all methods operate under the same reference–generation setting without manual selection bias. Specifically, we include the following models:
Text-based style-aligned image generation models.
We also compare our method against text-based style-aligned image generation models that leverage Stable Diffusion XL [podell2023sdxlimprovinglatentdiffusion] or FLUX [zhang2025alignedgen] as their backbone. Specifically, we include the following models:
Appendix B Comprehensive Analysis of our method
B.1 Additional analysis on generation process
To verify whether our method operates as intended during the generation process, we conduct an additional analysis at each step using 400 images generated from our method and the baseline model, Infinity [han2024infinity]. Specifically, we compare the first image in each batch with the subsequent images at different generation steps to examine how RGB statistics, style, and content evolve over time.